# On obstructions to the Euler system method for Rankin-Selberg convolutions

### v3

@@ Introduction
    $k,l\geq 2$,
@@
[…]
loefflerqnspecialQuestion 1  (Loeffler, see [15, Remark 4.4.2]) Assume that $k\geq 2$, $l=1$, $f$ does not have complex multiplication, and $\varepsilon_{f}\varepsilon_{g}\neq 1$. Is it true that, for all but finitely many $\mathfrak{p}$, there exists $\sigma\in G_{{\mathbb{Q}}(\mu_{p^{\infty}})}$ such that $T_{f,g,\mathfrak{p}}/(\sigma-1)T_{f,g,\mathfrak{p}}$ is free of rank one?
@@
[…]
weakercondsspecialTheorem A  Let $\mathcal{C}$ denote the collection of primitive Dirichlet characters $\chi$ such that $f$ has an inner twist by $\chi$. The answer to Question Introduction is positive when any of the following conditions holds:
@@
[…]
twistedbsdspecialTheorem B  Let $E$ be an elliptic curve over $\mathbb{Q}$} without complex multiplication and $\rho$ be an odd two-dimensional irreducible representation of $G_{{\mathbb{Q}}}$, with splitting field $K$ and coefficients in some number field $L_{0}$ with ring of integers $\mathcal{O}$. Let $\mathfrak{p}$ be a maximal ideal of $\mathcal{O}$ with residue characteristic $p$. Suppose that the following assumptions hold:
@@
[…]
weakerqnspecialTheorem C  If one replaces “for all but finitely many $\mathfrak{p}$” with “for infinitely many $\mathfrak{p}$” in Question Introduction, then the answer is positive.
@@
[…]
inf-cexspecialTheorem D  (See Section 3) Let $f\in\mathcal{S}_{k}(\Gamma_{1}(N))$ be a newform with character $\varepsilon_{f}$ without complex multiplication. Suppose that one of the following conditions holds:
@@
[…]
counterexspecialTheorem E  (See Proposition 4 (d)) There exist newforms $f\in\mathcal{S}_{2}(\Gamma_{0}(63)),g\in\mathcal{S}_{1}(\Gamma_{1}(1452))$ with coefficients respectively in ${\mathbb{Q}}(\sqrt{3}),{\mathbb{Q}}(\sqrt{-3})$ satisfying the following property: for all but finitely many prime ideals $\mathfrak{p}$ of $L={\mathbb{Q}}(e^{2i\pi/12})$ with residue characteristic $p\equiv 5,7\pmod{12}$, there is no $\sigma\in G_{{\mathbb{Q}}(\mu_{p^{\infty}})}$ such that $\ker\left(\sigma-\mathrm{id}\mid T_{f,\mathfrak{p}}\otimes_{\mathcal{O}_{L_{\mathfrak{p}}}}T_{g,\mathfrak{p}}\right)\otimes_{\mathcal{O}_{L_{\mathfrak{p}}}}L_{\mathfrak{p}}$ is a $L_{\mathfrak{p}}$-line.
@@
Then should one expect, as in [10, Theorem 11.7.3], that the Nekovár̆ cohomology group $\tilde{H}^{2}\left(\mathcal{O}_{L}\left[\frac{1}{N_{f}N_{g}p}\right],T_{f,\mathfrak{p}}\otimes T_{g,\mathfrak{p}};\Delta^{BK}\right)$ is finite [^2] The statement of the theorem in loc. cit. seems to contain a slight typo; one should (probably) read $M_{\mathcal{O}}(f\otimes g)^{\ast}$ instead of $M_{\mathcal{O}}(f\otimes g)$.? However, how to go about proving such a claim without the theory of Euler systems does not seem clear at all.
Acknowledgements   This question arose in the context of my PhD thesis, and I am grateful to my supervisor Loïc Merel for encouraging me to work on this project. His careful reading and many comments helped me vastly improve the original version of this text. I would like to thank Pierre Parent and Jan Vonk for their comments as well. I would also like to thank David Loeffler for his help regarding the application of the results of this paper to the twisted Birch and Swinnerton-Dyer conjecture. I would finally like to thank the anonymous reviewers for their attentive reading of this article and their many remarks, comments and suggestions.
@@ 1 The big image theorem for one modular form
[…]
twist-existsspecial Proposition 1.1  Let $f\in\mathcal{S}_{k}(\Gamma_{1}(N))$ be a newform with character $\varepsilon_{f}$ of conductor $M$. Let $\chi$ be a primitive Dirichlet character of conductor $C$. There exists a unique new, cuspidal eigenform $f\otimes\chi\in\mathcal{S}_{k}(\Gamma_{1}(L))$ such that, for every prime number $p\nmid NC$, $a_{p}(f\otimes\chi)=a_{p}(f)\chi(p)$. We say that $f\otimes\chi$ is the twist of $f$ by $\chi$. Moreover:
@@
[…]
    twist-characterspecial The character of $f\otimes\chi$ is the primitive Dirichlet character associated to $\varepsilon_{f}\chi^{2}$.
@@
[…]
    level-twist-basicspecial $L$ divides the least common multiple of $N,CM,C^{2}$.
@@
[…]
    level-twist-coprimespecial For every prime number $q\nmid C$, one has $v_{q}(L)=v_{q}(N)$.
@@
[…]
    level-twist-localspecial Let $q$ be a prime number, and write $\chi=\chi_{q}\chi^{\prime}$, where the conductor of $\chi_{q}$ (resp. $\chi^{\prime}$) is a power of $q$ (resp. coprime to $q$). Then $v_{q}(L)$ only depends on $f$ and $\chi_{q}$.
@@
Let $f\in\mathcal{S}_{k}(\Gamma_{1}(N))$ be a newform with (primitive) character $\varepsilon_{f}$. Let $f(\tau)=\sum_{n\geq 1}{a_{n}(f)q^{n}}$ be its $q$-expansion and let $L={\mathbb{Q}}(a_{n}(f)\mid n\geq 1)$.
@@
[…]
momose15special Lemma 1.3  If $f$ has an inner twist by $(\gamma,\chi)$, then one can write $\chi=\lambda\varepsilon_{f}^{t}$, where $t\in{\mathbb{Z}}$ and $\lambda$ is quadratic. Moreover, the conductor of $\chi$ divides $N$ and one has $\gamma(L)=L$.
@@
[…]
abelian-dichospecial Proposition 1.4  The formula $(\alpha,\chi)\cdot(\beta,\psi)=(\alpha\circ\beta,\chi\alpha(\psi))$ endows $\Gamma_{f}$ with an abelian group law. Moreover, the group homomorphism
@@
Remark  While we will not be considering it in the rest of the paper, let us briefly discuss the situation when $k\geq 2$ and $f$ has CM by an imaginary quadratic field $K$. The following claims can essentially be deduced from the contents of [21, §4] (using Serre’s theorem on rational abelian semi-simple representations [23, Chap. III., §3.1, Theorem and Remark 1]), and in particular the fact that $f$ is constructed from a Grössencharakter $\psi$ as in [21, pp. 34–35].
@@ 2 Special elements in the image of the convolution of two modular forms
[…]
jointimage-Hspecial Proposition 2.1  Let $\mathfrak{p}$ be a good prime with residue characteristic $p$. The image of $G_{H(\mu_{p^{\infty}})}$ in $\operatorname{Aut}(T_{f,\mathfrak{p}})\times\operatorname{Aut}(T_{g})$ is $\operatorname{SL}_{2}(\mathcal{O}_{F_{\mathfrak{p}}})\times\rho_{g}(G_{H})$. Moreover, for any $\sigma\in G_{{\mathbb{Q}}}$, there exists $\sigma^{\prime}\in G_{{\mathbb{Q}}(\mu_{p^{\infty}})}$ such that $\rho_{g}(\sigma)=\rho_{g}(\sigma^{\prime})$ and $\sigma,\sigma^{\prime}$ have the same image in $\mathrm{Gal}(H/{\mathbb{Q}})$.
@@
Recall that, for a Hausdorff topological group $G$, its derived subgroup is the closed subgroup $G^{\prime}$ generated by its commutators, i.e. its elements of the form $ghg^{-1}h^{-1}$ with $g,h\in G$. The subgroup $G^{\prime}$ is preserved by every continuous automorphism of $G$, and it is contained in the kernel of every continuous map $G\rightarrow A$ where $A$ is a Hausdorff topological abelian group.
@@
[…]
papierspecial Corollary 2.2  Let $\mathfrak{p}$ be a good prime with residue characteristic $p$ and $\sigma\in G_{{\mathbb{Q}}}$. There exists $\alpha\in\mathcal{O}_{L,\mathfrak{p}}^{\times}$ such that for all $(\gamma,\chi)\in\Gamma_{f}$, one has $\gamma(\alpha)=\chi(\sigma)\alpha$. If moreover $\sigma\in G_{{\mathbb{Q}}(\mu_{p^{\infty}})}$, the image of $\sigma G_{H(\mu_{p^{\infty}})}$ in $\operatorname{Aut}(T_{f,\mathfrak{p}})\times\operatorname{Aut}(T_{g})$ is $\left[\begin{pmatrix}\alpha&0\\ 0&\alpha^{-1}\varepsilon_{f}(\sigma)\end{pmatrix}\operatorname{SL}_{2}(\mathcal{O}_{F_{\mathfrak{p}}})\right]\times\left[\rho_{g}(\sigma)\rho_{g}(G_{H})\right]$.
@@
[…]
irred-product-abstractspecial Lemma 2.3  Let $V,W$ be finite-dimensional vector spaces over a field $k$, and $G_{1},G_{2}$ be subgroups of $\mathrm{GL}(V)$ and $\mathrm{GL}(W)$. Let $\Gamma$ be a subgroup of $\mathrm{GL}(V)\times\mathrm{GL}(W)$, containing $G_{1}\times\{\mathrm{id}\}$ and such that its second projection is $G_{2}$. Suppose that $V$ is an absolutely irreducible $k[G_{1}]$-module and $W$ is an irreducible $k[G_{2}]$-module. Then $V\otimes W$ is an irreducible $k[\Gamma]$-module.
@@
[…]
only-we-sespecial Proposition 2.4  If $\mathfrak{p}$ is a good prime of residue characteristic $p$, $T_{f,g,\mathfrak{p}}$ satisfies conditions (N) and (gI). Moreover, if $Q/2$ is not a power of $p$, then condition (rI) holds.
@@
Let $h\in K$. Then $h$ acts on $T_{g}/\mathfrak{p}$ with characteristic polynomial $(X-u)^{2}$, for some $u\in(\mathcal{O}_{L_{0}}/\mathfrak{p})^{\times}$ with Teichmüller lift $u_{0}\in\mathcal{O}_{(L_{0})_{\mathfrak{p}}}^{\times}$. Then $h^{\prime}:=u_{0}^{-1}h\in\operatorname{GL}_{2}(\mathcal{O}_{(L_{0})_{\mathfrak{p}}})$ has finite order and lies in a conjugate of the pro-$p$-group $\{M\in\operatorname{GL}_{2}(\mathcal{O}_{(L_{0})_{\mathfrak{p}}}),\,M\equiv\begin{pmatrix}1&\ast\\ 0&1\end{pmatrix}\pmod{\mathfrak{p}}\}$. Therefore, the order of $h^{\prime}$ is a power of $p$, so that $h^{p^{n}}$ is scalar for some $n\geq 1$.
@@
Remark  Using the fact that $\operatorname{SL}_{2}(\mathcal{O}_{F_{\mathfrak{p}}})$ is its own derived subgroup, we can slightly refine the proof of Proposition 2 to find $\tau\in G_{{\mathbb{Q}}}^{\prime}$ such that acts on $T_{f,g,\mathfrak{p}}$ by $\mathrm{id}$. Let now $\eta:G_{{\mathbb{Q}}}\rightarrow\{\pm 1\}$ be a quadratic character and assume that $T_{f,g,\mathfrak{p}}$ satisfies (wE) (resp. (sE)). Let $\sigma\in G_{{\mathbb{Q}}(\mu_{p^{\infty}})}$ be such that $\ker(\sigma-1\mid T_{f,g,\mathfrak{p}})\otimes(L_{0})_{\mathfrak{p}}$ is a line (resp. $T_{f,g,\mathfrak{p}}/(\sigma-1)T_{f,g,\mathfrak{p}}$ is free of rank one over $\mathcal{O}_{(L_{0})_{\mathfrak{p}}}$).
@@
The above Proposition is not sufficient in order to answer Question Introduction: in general, we cannot assume that we can find $\alpha,u$ that are both contained in $F_{\mathfrak{p}}$. There is however one special case in which the argument still applies: when $\alpha=1$ and $u=\pm 1$. The goal of the two following Propositions is to give sufficient conditions that imply that $\rho_{g}(G_{H})$ contains a non-scalar matrix with $\pm 1$ as an eigenvalue.
[…]
special2special Proposition 2.6  Assume that $1\in\varepsilon_{g}(G_{H})$ and that $\rho_{g}(G_{H})$ does not contain an element similar to $\Delta_{-1,1}$. Then $g$ has either real or complex multiplication by some quadratic field $K$. If moreover we can choose $K$ so that it is not contained in $H$, then the following holds:
@@
Proof. –  Let $M\in\operatorname{GL}_{2}(L_{0})$ be a matrix with $\det{M}=-1$ whose image in $\operatorname{PGL}_{2}(L_{0})$ has order $2$. Then $M^{2}$ is scalar, so $M$ is conjugate in $\operatorname{GL}_{2}({\mathbb{C}})$ to $\Delta_{a,b}$ for some $a,b\in{\mathbb{C}}^{\times}$ with a=ba=-b and ab=1ab=-1. Therefore $M$ is similar to $\Delta_{1,-1}$ in $\operatorname{GL}_{2}({\mathbb{C}})$, so $M$ is similar to $\Delta_{1,-1}$ in $\operatorname{GL}_{2}(L_{0})$.
@@
[…]
specialqspecial Proposition 2.7  Assume that there exists an odd prime $q\mid|\tilde{\rho}_{g}(G_{{\mathbb{Q}}})|$ such that $\varepsilon_{g}(G_{H})$ contains an element of order $q$. Then one of the following statements holds:
@@
    $q=3$, $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ and $\tilde{\rho}_{g}(G_{H})$ are isomorphic to $A_{4}$, $e^{2i\pi/9}\in\varepsilon_{g}(G_{H})$, and the projection $\rho_{g}(G_{H})\cap\operatorname{SL}_{2}(L_{0})\rightarrow\tilde{\rho}_{g}(G_{H})$ is not surjective.
@@
Because $H/{\mathbb{Q}}$ is abelian, one has $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})^{\prime}\subset\tilde{\rho}_{g}(G_{H})\subset\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$, so either $\tilde{\rho}_{g}(G_{H})=\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$, or $\tilde{\rho}_{g}(G_{H})=\tilde{\rho}_{g}(G_{{\mathbb{Q}}})^{\prime}\simeq({\mathbb{Z}}/2{\mathbb{Z}})^{\oplus 2}$. The result of the Proposition follows from the study of the following three cases and the argument at the beginning of the proof.
@@
[…]
sufficient-for-euler-adaptedspecial Corollary 2.8  Let $\mathfrak{p}$ be a good prime for $(f,g)$. If $\rho_{g}(G_{H})$ contains a non-scalar matrix $M$ such that $1$ or 1-1 is an eigenvalue of $M$, then $T_{f,g,\mathfrak{p}}$ satisfies (wE); if furthermore the order of $M$ is prime to $p$, then $T_{f,g,\mathfrak{p}}$ satisfies (sE). Hence $T_{f,g,\mathfrak{p}}$ has Euler-adapted Galois image for all but finitely many primes $\mathfrak{p}$. In particular, all of the above applies if any of the following conditions is satisfied:
@@
[…]
    sc1special $\rho_{g}(G_{H})$ contains the image under $\rho_{g}$ of the complex conjugation. This is the case if $H$ is unramified at every prime dividing the level of $g$.
@@
[…]
    sc2special $1\in\varepsilon_{g}(G_{H})$ and $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is not dihedral.
@@
[…]
    sc3special $1\in\varepsilon_{g}(G_{H})$, $\tilde{\rho}_{g}(G_{H})$ is a dihedral group of order $4n$ with $n\geq 1$.
@@
[…]
    sc4special $1\in\varepsilon_{g}(G_{H})$, $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is a dihedral group of order $8n$ with $n\geq 1$, and $g$ has either real or complex multiplication by some quadratic field $K\not\subset H$.
@@
[…]
    sc5special $\varepsilon_{g}(G_{H})$ contains 1-1 but not $i$, and, for some quadratic number field $K$ by which $g$ has either real or complex multiplication, one has $K\not\subset H$.

### v4

@@ Introduction
    $k,l\geq 2$ and $k$ or $l$ is odd [^1] Loeffler’s results apply somewhat more generally, but this is a simple sufficient condition.,
@@
Question 1  (Loeffler, see [15, Remark 4.4.2]) Assume that $k\geq 2$, $l=1$, $f$ does not have complex multiplication, and $\varepsilon_{f}\varepsilon_{g}\neq 1$. Is it true that, for all but finitely many $\mathfrak{p}$, there exists $\sigma\in G_{{\mathbb{Q}}(\mu_{p^{\infty}})}$ such that $T_{f,g,\mathfrak{p}}/(\sigma-1)T_{f,g,\mathfrak{p}}$ is free of rank one?
@@
Theorem A  Let $\mathcal{C}$ denote the collection of primitive Dirichlet characters $\chi$ such that $f$ has an inner twist by $\chi$. The answer to Question Introduction is positive when any of the following conditions holds:
@@
Theorem B  Let $E$ be an elliptic curve over $\mathbb{Q}$} without complex multiplication and $\rho$ be an odd two-dimensional irreducible representation of $G_{{\mathbb{Q}}}$, with splitting field $K$ and coefficients in some number field $L_{0}$ with ring of integers $\mathcal{O}$. Let $\mathfrak{p}$ be a maximal ideal of $\mathcal{O}$ with residue characteristic $p$. Suppose that the following assumptions hold:
@@
Theorem C  If one replaces “for all but finitely many $\mathfrak{p}$” with “for infinitely many $\mathfrak{p}$” in Question Introduction, then the answer is positive.
@@
Theorem D  (See Section 3) Let $f\in\mathcal{S}_{k}(\Gamma_{1}(N))$ be a newform with character $\varepsilon_{f}$ without complex multiplication. Suppose that one of the following conditions holds:
@@
Theorem E  (See Proposition 4 (d)) There exist newforms $f\in\mathcal{S}_{2}(\Gamma_{0}(63)),g\in\mathcal{S}_{1}(\Gamma_{1}(1452))$ with coefficients respectively in ${\mathbb{Q}}(\sqrt{3}),{\mathbb{Q}}(\sqrt{-3})$ satisfying the following property: for all but finitely many prime ideals $\mathfrak{p}$ of $L={\mathbb{Q}}(e^{2i\pi/12})$ with residue characteristic $p\equiv 5,7\pmod{12}$, there is no $\sigma\in G_{{\mathbb{Q}}(\mu_{p^{\infty}})}$ such that $\ker\left(\sigma-\mathrm{id}\mid T_{f,\mathfrak{p}}\otimes_{\mathcal{O}_{L_{\mathfrak{p}}}}T_{g,\mathfrak{p}}\right)\otimes_{\mathcal{O}_{L_{\mathfrak{p}}}}L_{\mathfrak{p}}$ is a $L_{\mathfrak{p}}$-line.
@@
Then should one expect, as in [10, Theorem 11.7.3], that the Nekovár̆ cohomology group $\tilde{H}^{2}\left(\mathcal{O}_{L}\left[\frac{1}{N_{f}N_{g}p}\right],T_{f,\mathfrak{p}}\otimes T_{g,\mathfrak{p}};\Delta^{BK}\right)$ is finite [^3] The statement of the theorem in loc. cit. seems to contain a slight typo; one should (probably) read $M_{\mathcal{O}}(f\otimes g)^{\ast}$ instead of $M_{\mathcal{O}}(f\otimes g)$.? How to go about proving such a claim without the theory of Euler systems does not seem clear at all.
Acknowledgements   This question arose in the context of my PhD thesis, and I am grateful to my supervisor Loïc Merel for encouraging me to work on this project. His careful reading and many comments helped me vastly improve the original version of this text. I would like to thank Pierre Parent and Jan Vonk for their comments as well. I would also like to thank David Loeffler for his comments and his help regarding the application of the results of this paper to the twisted Birch and Swinnerton-Dyer conjecture. I would finally like to thank the anonymous reviewers for their attentive reading of this article and their many remarks, comments and suggestions. This research was partially funded by the European Union, ERC Starting Grant 101076941 ’GAGARIN’.
@@ 1 The big image theorem for one modular form
Proposition 1.1  Let $f\in\mathcal{S}_{k}(\Gamma_{1}(N))$ be a newform with character $\varepsilon_{f}$ of conductor $M$. Let $\chi$ be a primitive Dirichlet character of conductor $C$. There exists a unique new, cuspidal eigenform $f\otimes\chi\in\mathcal{S}_{k}(\Gamma_{1}(L))$ such that, for every prime number $p\nmid NC$, $a_{p}(f\otimes\chi)=a_{p}(f)\chi(p)$. We say that $f\otimes\chi$ is the twist of $f$ by $\chi$. Moreover:
@@
    The character of $f\otimes\chi$ is the primitive Dirichlet character associated to $\varepsilon_{f}\chi^{2}$.
@@
    $L$ divides the least common multiple of $N,CM,C^{2}$.
@@
    For every prime number $q\nmid C$, one has $v_{q}(L)=v_{q}(N)$.
@@
    Let $q$ be a prime number, and write $\chi=\chi_{q}\chi^{\prime}$, where the conductor of $\chi_{q}$ (resp. $\chi^{\prime}$) is a power of $q$ (resp. coprime to $q$). Then $v_{q}(L)$ only depends on $f$ and $\chi_{q}$.
@@
Let $f\in\mathcal{S}_{k}(\Gamma_{1}(N))$ be a newform with (primitive) character $\varepsilon_{f}$. Let $f(\tau)=\sum_{n\geq 1}{a_{n}(f)q^{n}}$ be its $q$-expansion and let $L={\mathbb{Q}}(a_{n}(f)\mid n\geq 1)\subset{\mathbb{C}}$.
@@
Lemma 1.3  If $f$ has an inner twist by $(\gamma,\chi)$, then one can write $\chi=\lambda\varepsilon_{f}^{t}$, where $t\in{\mathbb{Z}}$ and $\lambda$ is quadratic. Moreover, the conductor of $\chi$ divides $N$ and one has $\gamma(L)=L$.
@@
Proposition 1.4  The formula $(\alpha,\chi)\cdot(\beta,\psi)=(\alpha\circ\beta,\chi\alpha(\psi))$ endows $\Gamma_{f}$ with an abelian group law. Moreover, the group homomorphism
@@
Remark  While we will not be considering it in the rest of the paper, let us briefly discuss the situation when $k\geq 2$ and $f$ has CM by an imaginary quadratic field $K$. The following claims can essentially be deduced from the contents of [21, §4] (using Serre’s theorem on rational abelian semi-simple representations [23, Chap. III, §3.1, Theorem and Remark 1]), and in particular the fact that $f$ is constructed from a Grössencharakter $\psi$ as in [21, pp. 34–35].
@@ 2 Special elements in the image of the convolution of two modular forms
Proposition 2.1  Let $\mathfrak{p}$ be a good prime with residue characteristic $p$. The image of $G_{H(\mu_{p^{\infty}})}$ in $\operatorname{Aut}(T_{f,\mathfrak{p}})\times\operatorname{Aut}(T_{g})$ is $\operatorname{SL}_{2}(\mathcal{O}_{F_{\mathfrak{p}}})\times\rho_{g}(G_{H})$. Moreover, for any $\sigma\in G_{{\mathbb{Q}}}$, there exists $\sigma^{\prime}\in G_{{\mathbb{Q}}(\mu_{p^{\infty}})}$ such that $\rho_{g}(\sigma)=\rho_{g}(\sigma^{\prime})$ and $\sigma,\sigma^{\prime}$ have the same image in $\mathrm{Gal}(H/{\mathbb{Q}})$.
@@
Recall that, for a Hausdorff topological group $G$, its derived subgroup is the closed subgroup $G^{\prime}$ generated by its commutators, i.e. the elements of the form $ghg^{-1}h^{-1}$ with $g,h\in G$. The subgroup $G^{\prime}$ is preserved by every continuous automorphism of $G$, and it is contained in the kernel of every continuous map $G\rightarrow A$ where $A$ is a Hausdorff topological abelian group.
@@
Corollary 2.2  Let $\mathfrak{p}$ be a good prime with residue characteristic $p$ and $\sigma\in G_{{\mathbb{Q}}}$. There exists $\alpha\in\mathcal{O}_{L,\mathfrak{p}}^{\times}$ such that for all $(\gamma,\chi)\in\Gamma_{f}$, one has $\gamma(\alpha)=\chi(\sigma)\alpha$. If moreover $\sigma\in G_{{\mathbb{Q}}(\mu_{p^{\infty}})}$, the image of $\sigma G_{H(\mu_{p^{\infty}})}$ in $\operatorname{Aut}(T_{f,\mathfrak{p}})\times\operatorname{Aut}(T_{g})$ is $\left[\begin{pmatrix}\alpha&0\\ 0&\alpha^{-1}\varepsilon_{f}(\sigma)\end{pmatrix}\operatorname{SL}_{2}(\mathcal{O}_{F_{\mathfrak{p}}})\right]\times\left[\rho_{g}(\sigma)\rho_{g}(G_{H})\right]$.
@@
Lemma 2.3  Let $V,W$ be finite-dimensional vector spaces over a field $k$, and $G_{1},G_{2}$ be subgroups of $\mathrm{GL}(V)$ and $\mathrm{GL}(W)$. Let $\Gamma$ be a subgroup of $\mathrm{GL}(V)\times\mathrm{GL}(W)$, containing $G_{1}\times\{\mathrm{id}\}$ and such that its second projection is $G_{2}$. Suppose that $V$ is an absolutely irreducible $k[G_{1}]$-module and $W$ is an irreducible $k[G_{2}]$-module. Then $V\otimes W$ is an irreducible $k[\Gamma]$-module.
@@
Proposition 2.4  If $\mathfrak{p}$ is a good prime of residue characteristic $p$, $T_{f,g,\mathfrak{p}}$ satisfies conditions (N) and (gI). Moreover, if $Q/2$ is not a power of $p$, then condition (rI) holds.
@@
Let $h\in K$. Then $h$ acts on $T_{g}/\mathfrak{p}$ with characteristic polynomial $(X-u)^{2}$, for some $u\in(\mathcal{O}_{L_{0}}/\mathfrak{p})^{\times}$ with Teichmüller lift $u_{0}\in\mathcal{O}_{(L_{0})_{\mathfrak{p}}}^{\times}$. Then $h^{\prime}:=u_{0}^{-1}h\in\operatorname{GL}_{2}(\mathcal{O}_{(L_{0})_{\mathfrak{p}}})$ has finite order and its characteristic polynomial is congruent modulo $\mathfrak{p}$ to $(X-1)^{2}$. Therefore, the order of $h^{\prime}$ is a power of $p$, so that $h^{p^{n}}$ is scalar for some $n\geq 1$.
@@
Remark  Using the fact that $\operatorname{SL}_{2}(\mathcal{O}_{F_{\mathfrak{p}}})$ is its own derived subgroup, we can slightly refine the proof of Proposition 2 to find $\tau\in G_{{\mathbb{Q}}}^{\prime}$ acting on $T_{f,g,\mathfrak{p}}$ by $\mathrm{id}$. Let now $\eta:G_{{\mathbb{Q}}}\rightarrow\{\pm 1\}$ be a quadratic character and assume that $T_{f,g,\mathfrak{p}}$ satisfies (wE) (resp. (sE)). Let $\sigma\in G_{{\mathbb{Q}}(\mu_{p^{\infty}})}$ be such that $\ker(\sigma-1\mid T_{f,g,\mathfrak{p}})\otimes(L_{0})_{\mathfrak{p}}$ is a line (resp. $T_{f,g,\mathfrak{p}}/(\sigma-1)T_{f,g,\mathfrak{p}}$ is free of rank one over $\mathcal{O}_{(L_{0})_{\mathfrak{p}}}$).
@@
Proposition 2 is not sufficient in order to answer Question Introduction: in general, we cannot assume that we can find $\alpha,u$ that are both contained in $F_{\mathfrak{p}}$. There is however one special case in which the argument still applies: when $\alpha=1$ and $u=\pm 1$. The goal of the two following Propositions is to give sufficient conditions that imply that $\rho_{g}(G_{H})$ contains a non-scalar matrix with $\pm 1$ as an eigenvalue.
Proposition 2.6  Assume that $1\in\varepsilon_{g}(G_{H})$ and that $\rho_{g}(G_{H})$ does not contain an element similar to $\Delta_{-1,1}$. Then $g$ has either real or complex multiplication by some quadratic field $K$. If moreover we can choose $K$ so that it is not contained in $H$, then the following holds:
@@
Proof. –  Let $M\in\operatorname{GL}_{2}(L_{0})$ be a matrix with $\det{M}=-1$ and whose image in $\operatorname{PGL}_{2}(L_{0})$ has order $2$. Then $M^{2}$ is scalar, so $M$ is conjugate in $\operatorname{GL}_{2}({\mathbb{C}})$ to $\Delta_{a,b}$ for some $a,b\in{\mathbb{C}}^{\times}$ with a=ba=-b and ab=1ab=-1. Therefore $M$ is similar to $\Delta_{1,-1}$ in $\operatorname{GL}_{2}({\mathbb{C}})$, so $M$ is similar to $\Delta_{1,-1}$ in $\operatorname{GL}_{2}(L_{0})$.
@@
Proposition 2.7  Assume that there exists an odd prime $q\mid|\tilde{\rho}_{g}(G_{{\mathbb{Q}}})|$ such that $\varepsilon_{g}(G_{H})$ contains an element of order $q$. Then one of the following statements holds:
@@
    $q=3$, $\tilde{\rho}_{g}(G_{H})=\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is isomorphic to $A_{4}$, $e^{2i\pi/9}\in\varepsilon_{g}(G_{H})$, and the projection $\rho_{g}(G_{H})\cap\operatorname{SL}_{2}(L_{0})\rightarrow\tilde{\rho}_{g}(G_{H})$ is not surjective.
@@
Because $H/{\mathbb{Q}}$ is abelian, one has $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})^{\prime}\subset\tilde{\rho}_{g}(G_{H})\subset\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$, so either $\tilde{\rho}_{g}(G_{H})=\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$, or $\tilde{\rho}_{g}(G_{H})=\tilde{\rho}_{g}(G_{{\mathbb{Q}}})^{\prime}\simeq({\mathbb{Z}}/2{\mathbb{Z}})^{\oplus 2}$. By the argument at the beginning of the proof, it is enough to treat the following three cases.
@@
Corollary 2.8  Let $\mathfrak{p}$ be a good prime for $(f,g)$. If $\rho_{g}(G_{H})$ contains a non-scalar matrix $M$ such that $1$ or 1-1 is an eigenvalue of $M$, then $T_{f,g,\mathfrak{p}}$ satisfies (wE); if furthermore the order of $M$ is prime to $p$, then $T_{f,g,\mathfrak{p}}$ satisfies (sE). Hence $T_{f,g,\mathfrak{p}}$ has Euler-adapted Galois image for all but finitely many primes $\mathfrak{p}$. In particular, all of the above applies if any of the following conditions is satisfied:
@@
    $\rho_{g}(G_{H})$ contains the image under $\rho_{g}$ of the complex conjugation. This is the case if $H$ is unramified at every prime dividing the level of $g$.
@@
    $1\in\varepsilon_{g}(G_{H})$ and $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is not dihedral.
@@
    $1\in\varepsilon_{g}(G_{H})$, $\tilde{\rho}_{g}(G_{H})$ is a dihedral group of order $4n$ with $n\geq 1$.
@@
    $1\in\varepsilon_{g}(G_{H})$, $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is a dihedral group of order $8n$ with $n\geq 1$, and $g$ has either real or complex multiplication by some quadratic field $K\not\subset H$.
@@
    $\varepsilon_{g}(G_{H})$ contains 1-1 but not $i$, and, for some quadratic number field $K$ by which $g$ has either real or complex multiplication, one has $K\not\subset H$.